A student randomly guesses the answers to a five question true/false test. If there is a 50% chance of guessing correctly on each question, what is the probability that the student misses exactly 1 question?
Answer by stanbon(75887)

A student randomly guesses the answers to a five question true/false test. If there is a 50% chance of guessing correctly on each question, what is the probability that the student misses exactly 1 question?
-----------------
It's a binomial problem with n=5, p=1/2, x=1
-------------------------------
P(x=1) = 5C1(1/2)^1(1/2)^4 = 5*1/32 = 5/32
=============================================
Cheers,
Stan H.